the fan has total head of 190 m and A static pressure of 20 cm wg. If the air density is 1.15 kg/m3, what is the velocity of air flowing.


1
Expert's answer
2021-10-08T00:53:07-0400

"\rho_{air}*h_a=\rho_{water}*h_w"

"h_a=\frac{200*10^{3}*10^{-3}}{1.15}"

=173.913

"\u2206h=190-173.913"

"=16.087"

"\frac{P_1-P_2}{\rho g}=\frac{V_2^{2}-V_1^{2}}{2g}"

"\frac{P_1-P_2}{\rho}=\frac{V_2^{2}-V_1^{2}}{2}"

V1=0

"\therefore V_2=\sqrt{2g(∆h)}"

"=\sqrt{2*9.81*16.087}"

"=\sqrt{315.62694}"

"=17.7659m\/s"